# Akash Srivastava2,3, Hongdong Li9, Junzhou Huang10, & Dimitris N. Metaxas1

### v2

@@ Abstract
Recent advances in discrete diffusion models have demonstrated strong performance in image generation and masked language modeling, yet they remain limited in their capacity for controlled content editing. We propose DICE (Discrete Inversion for Controllable Editing), a novel framework that pioneers precise inversion capabilities for discrete diffusion models, including both masked generative and multinomial diffusion variants. Our key innovation lies in capturing noise sequences and masking patterns during reverse diffusion process, enabling both accurate reconstruction and flexible editing without relying on predefined masks or attention-based manipulations. Through comprehensive experiments across image and text modalities using models such as Paella, VQ-Diffusion, and RoBERTa, we demonstrate that DICE successfully maintains high fidelity to the original data while significantly expanding editing capabilities. These results establish new possibilities for fine-grained content manipulation in discrete spaces.
@@ 1 Introduction
Diffusion models have emerged as a powerful class of generative models, achieving remarkable success in high-fidelity image and video synthesis [23, 12, 45, 42, 24, 39, 40, 10]. These models generate data by iteratively denoising samples from a simple noise distribution, effectively reversing a diffusion process that gradually corrupts data. Broadly, diffusion models can be categorized into continuous and discrete types, each tailored to different data modalities.
Figure 1: Illustration of the limitation of masked inpainting method. Here, we want to change the cat to a dog. Inpainting with masked generation inadvertently modifies the orientation of the head, resulting in a less favourable result. With our discrete inversion method, we are able to edit the image while preserving other properties of the object being edited. This is achieved by injecting the information from the input image into the logit space. Dotted red box indicates the mask.
Continuous diffusion models operate in continuous spaces, leveraging stochastic differential equations (SDEs) or their deterministic counterparts, ordinary differential equations (ODEs), to model the forward and reverse diffusion processes [49, 47]. Advances such as flow matching [29, 30, 9] have enhanced their efficiency and flexibility. These models have been successfully applied in various domains, including image editing [35, 3, 38, 18, 19, 59], medical imaging [20], and solving inverse problems [8, 50]. In image editing, continuous diffusion models enable controlled manipulation of images while preserving consistency with the underlying data distribution. A key capability enabling this is inversion—the process of reversing the diffusion model to recover the original noise vector or latent representation that could have generated a given data sample. Two main inversion approaches exist: deterministic inversion using ODEs (e.g., DDIM Inversion [47]) and stochastic inversion by recording noise sequences (e.g., CycleDiffusion [56], DDPM Inversion [12]).
Discrete diffusion models are designed for inherently discrete data such as text or image tokens [14]. They adapt the diffusion framework to discrete spaces by defining appropriate transition kernels that corrupt and restore discrete data [25, 2, 16]. Prominent examples include multinomial diffusion [25, 16], D3PM [2], and masked generative models like MaskGIT [5], Muse [6]. Despite their success in generation tasks, discrete diffusion models face limitations in controlled content editing. For instance, masked generative models achieve image editing through masked inpainting, where regions are masked and regenerated based on new conditions. However, this approach lacks the ability to inject information from the masked area into the inpainting process, limiting fine-grained control over the editing outcome, as illustrated in Figure 1.
@@
Moreover, existing ODE-based inversion techniques developed for continuous diffusion models are not directly applicable to discrete diffusion models due to inherent differences in data representation and diffusion processes. This gap hinders the ability to perform precise inversion and controlled editing in discrete spaces. To address this challenge, we propose DICE (Discrete Inversion for Controllable Editing), the first inversion algorithm for discrete diffusion models to the best of our knowledge. Our method extends the stochastic inversion approach to discrete diffusion models, including both multinomial diffusion and masked generative models. The core idea is to record the noise sequence needed to recover a stochastic trajectory in the reverse diffusion process. Specifically, given an artificial trajectory where latent states have low correlation, we fit reverse sampling steps to this trajectory and save the residuals between targets and predictions. This process imprints the information of the original input data into the recorded residuals. During editing or inference, the residuals are added back, allowing us to inject and control the amount of information introduced into the inference process.
Our approach enables accurate reconstruction of the original input data and facilitates controlled editing without the need for predefined masks or attention map manipulation. It provides a flexible framework for fine-grained content manipulation in discrete spaces, overcoming the limitations of existing methods. We validate the effectiveness of DICE through extensive experiments on both image and text modalities. We evaluate our method on models such as VQ-Diffusion [16], Paella [43], and RoBERTa [31], demonstrating its versatility across different types of discrete generative models. Additionally, we introduce a novel text-editing dataset to further showcase our method’s capabilities and to facilitate future research in this area. Our contributions can be summarized as follows:
@@ 2 Related Work
Discrete diffusion. Generative modeling over discrete spaces using diffusion models, analogous to continuous space diffusion, was first introduced in [46]. Argmax flows [25] and D3PM [2] further developed this concept by modeling the noising process as a discrete-time, discrete-space Markov chain. These approaches utilize a variational loss function to train a neural network to reverse the Markov chain. Following those works, [13] and [16] introduced the VQ-GAN as a way to discretize the image into tokens. This enables efficient non-autoregressive methods, including MaskGIT [5], Muse [6] and MMVID [17], for image generation by iteratively remasking and reprediction. Building on the continuous counterpart, which is trained via score matching as a regression task, recent work [32, 36] introduces a discrete analogue, termed ratio matching. This approach extends score matching [48] to discrete spaces by learning probability ratios. Continuing this trend, [15] proposed discrete flow matching. In natural language processing, BERT [11] and RoBERTa [31] can also be viewed through the lens of discrete diffusion [52], using a similar discrete sequence modeling approach.
Diffusion inversion. Diffusion inversion aims to find an encoding or latent representation of the input signal that can be used to reconstruct the original data. Traditional approaches to diffusion inversion are based on neural ODEs [7], such as DDIM inversion [47] and flow matching [29, 30], where deterministic trajectories are used for inversion. Another class of methods focuses on stochastic differential equations (SDEs) [49], including models like CycleDiffusion [56] and DDPM Inversion [26], which rely on tracking noise or residuals along a stochastic path to recover the input. Our approach generalizes the concept of DDPM Inversion by extending it to discrete diffusion models, enabling effective inversion in both continuous and discrete settings.
Inversion-based image editing. DDIM inversion [47] has served as a foundational technique for various diffusion-based image editing approaches. In many image editing tasks, DDIM-type methods are often employed alongside guidance techniques like Prompt-to-Prompt [21], which manipulate cross-attention maps, as well as self-attention maps, as demonstrated by approaches like Plug-and-Play [51], TF-ICON [33], and StyleAligned [22]. On the other hand, DDPM inversion-based approaches [26] are known for their user-friendly nature, as they typically do not require complex attention map manipulations. These approaches are also versatile and can integrate with semantic guidance techniques, such as SEGA and LEDITS++, enabling broader applicability. To address challenges such as inaccurate reconstruction and error accumulation, Null-text Inversion [38] introduces test-time optimization of null embeddings, ensuring the reconstruction trajectory aligns more closely with the DDIM inversion path. Negative-prompt Inversion [37, 19] further improves time efficiency by providing a closed-form solution to an approximate inversion problem, reducing computational costs while maintaining competitive reconstruction quality.
@@ 3.1 Preliminaries
Masked generative modeling. Masked generative modeling is widely used in representation learning for both natural language processing and computer vision. It works by masking parts of the input and training the model to reconstruct the missing data. In models like BERT [11] and RoBERTa [31], masked tokens ([MASK]) are predicted based on the surrounding context, excelling in text completion and embedding representation learning. For image generation, Paella [43] adapts this approach for text-conditional image generation by renoising tokens instead of masking. The inference process in masked generative models typically involves iterative renoise/remask and repredict steps.
Multinomial Diffusion. Denoting $\bm{x}_{0}\in\{1,\ldots,K\}^{D}$ as a data point of dimension $D$. We use $\bm{v}({x}_{t}^{(i)})$ to denote the one hot column vector representation of the $i$-th entry of $\bm{x}_{t}$. To simplify notation, in the following we drop index $i$ and any function that operates on vector $\bm{x}_{t}$ is populated along its dimension. Diffusion model defines a markov chain $q(\bm{x}_{1:T}|\bm{x}_{0})=\Pi_{t=1}^{T}q(\bm{x}_{t}|\bm{x}_{t-1})$ that gradually add noise to the data $\bm{x}_{0}$ for $T$ times so that $\bm{x}_{T}$ contains little to no information. Discrete diffusion model [25, 2, 16] proposed an alternative likelihood-based model for categorical data, and defines the forward process following:
@@
with $p_{\theta}(\tilde{x}_{0}|x_{t})$ is parameterized by a neural network. We gradually denoise from $x_{T}$ to $x_{0}$ using 3. For numerical stability, the implementation uses log space instead of probability space. Masked generative models can be viewed as a special case of multinomial diffusion models with an additional absorbing state (or the [MASK] state). Its training objective can be viewed as a reweighted ELBO [4].
3.2 Discrete Inversion for Controllable Editing
Non ODE-based inversion. ODE-based generative models, such as DDIM and flow matching, define an ODE trajectory. Due to the deterministic nature of ODEs, inversion can be achieved by solving the ODE using the Euler method in forward direction, ensuring reconstruction based on the inherent properties of the ODE. In contrast, another line of research focuses on SDE-based models, such as CycleDiffusion [56] and DDPM Inversion [26]. Broadly speaking, these approaches ensure reconstruction by recording the noises or residuals that are required to reproduce the stochastic trajectory. CycleDiffusion records the Gaussian noise $\bm{z}_{t}$ during sampling from posterior $p(\bm{x}_{t-1}|\bm{x}_{t},\bm{x}_{0}=\bm{x}_{0})$ and injects information of the input signal by feeding the true $\bm{x}_{0}$. DDPM Inversion, on the other hand, incorporates information into $\bm{z}_{t}$ by fitting the reverse process into an artificial stochastic trajectory obtained by independent q-sample. For both CycleDiffusion and DDPM Inversion, the key idea is to utilize the Gaussian reparameterization trick, $x=\mu+\sigma z\Leftrightarrow x\sim\mathcal{N}(x;\mu,\sigma^{2})$, and keeping track of the “noise” that could have generated the sample from mean. For discrete diffusion models, we utilize the Gumbel-Max trick [34, 27], $x=\operatorname*{arg\,max}{\left(\log(\bm{\pi})+\bm{g}\right)}\Leftrightarrow x\sim\text{Cat}(x;\bm{\pi})$. Figure 2 provides an intuition of the proposed method.
@@ 3.2 Discrete Inversion for Controllable Editing
16:     $\bm{x}_{t-1}\leftarrow\tilde{\bm{x}}_{0}\odot(\bm{1}-\bm{m}_{t-1})+\bm{n}\odot\bm{m}_{t-1}$triangleright Re-noise
@@
2:for $t$ from $1$ to $T$ do
3:     $\bm{x}_{t}\sim q(\bm{x}_{t}|\bm{x}_{0})$triangleright Independent q-sample using Eq 5
@@
6:for $t$ from $T$ to $1$ do
7:     $\hat{\bm{y}}_{t-1}\leftarrow\log(\bm{\pi}_{\theta}(\bm{x}_{t},\bm{c},t))$triangleright Log posterior using Eq 3
8:     $\bm{z}_{t}\leftarrow\bm{y}_{t-1}-\hat{\bm{y}}_{t-1}$triangleright Eq 6
@@
Method PSNR $\uparrow$ LPIPS×${}_{{}^{\times 10^{3}}}$downarrow MSE×${}_{{}^{\times 10^{4}}}$downarrow SSIM×${}_{{}^{\times 10^{2}}}$uparrow
Inpainting 10.50 565.11 1002.09 30.13
Ours 30.91 39.81 11.07 90.22
$\text{Ours}^{\dagger}$ Inf 0.07 0.01 99.99
Table 1: Inversion Reconstruction performance. $\dagger$ The metric is calculated between the original image and its inverted counterpart. Due to the encoding and decoding steps in the VQ-VAE/GAN process, some inaccuracies are introduced by the quantization. The PSNR is Inf due to the reconstruction of our method yielding the same VQ-VAE/GAN latents. Base model is Paella [43].
Inverting masked generative models. In masked generative modeling, the stochastic trajectory ${\bm{x}_{t}}$ is constructed according to the specific inference algorithm of the model in use. For example, in Paella [43], the masking is inclusive, meaning that as the time step $t$ increases, the set of masked tokens grows. In contrast, the Unleashing Transformer [4] employs random masking at each step, where masks are generated independently using the q-sample function. Without loss of generality, we define a denoiser function $\mathcal{D}_{\theta}$ (parameterized by $\theta$). This denoiser outputs the logits of the predicted unmasked data given the noisy tokens $\bm{x}_{t}$. Unlike DDPM or multinomial diffusion, where $x_{t-1}$ is not sampled from a posterior given $x_{t}$, the inference of masked modeling takes a different approach. In masked modeling, $x_{t}$ is obtained from sampled $\hat{x}_{0|t}$ by re-noising. Since the categorical sampling happens at sampling from the denoiser’s prediction, we therefore define an corresponding latent sequence:

### v3

@@ Abstract
Recent advances in discrete diffusion models have demonstrated strong performance in image generation and masked language modeling, yet they remain limited in their capacity for controlled content editing. We propose DICE (Discrete Inversion for Controllable Editing), a novel framework that pioneers precise inversion capabilities for discrete diffusion models, including both masked generative and multinomial diffusion variants. Our key innovation lies in capturing noise sequences and masking patterns during reverse diffusion process, enabling both accurate reconstruction and flexible editing without relying on predefined masks or attention-based manipulations. Through comprehensive experiments across image and text modalities using models such as Paella, VQ-Diffusion, RoBERTa and LLaDA, we demonstrate that DICE successfully maintains high fidelity to the original data while significantly expanding editing capabilities. These results establish new possibilities for fine-grained content manipulation in discrete spaces.
@@ 1 Introduction
Continuous diffusion models operate in continuous spaces, leveraging stochastic differential equations (SDEs) or their deterministic counterparts, ordinary differential equations (ODEs), to model the forward and reverse diffusion processes (song2020score, ; song2021denoising, ).
Advances such as flow matching (lipman2022flow, ; liu2022flow, ; dao2023flow, ) have enhanced their efficiency and flexibility. These models have been successfully applied in various domains, including image editing (meng2021sdedit, ; avrahami2022blended, ; mokady2022null, ; han2023svdiff, ; han2024proxedit, ; zhang2023sine, ), medical imaging (he2023dmcvr, ), and solving inverse problems (chung2022diffusion, ; stathopoulos2024score, ). In image editing, continuous diffusion models enable controlled manipulation of images while preserving consistency with the underlying data distribution. A key capability enabling this is inversion—the process of reversing the diffusion model to recover the original noise vector or latent representation that could have generated a given data sample. Two main inversion approaches exist: deterministic inversion using ODEs (e.g., DDIM Inversion (song2021denoising, )) and stochastic inversion by recording noise sequences (e.g., CycleDiffusion (wu2022unifying, ), DDPM Inversion (huberman2024edit, )).
Figure 1: Illustration of the limitation of masked inpainting method. Inpainting with masked generation inadvertently modifies the orientation of the head, resulting in a less favourable result. With our discrete inversion method, we are able to edit the image while preserving other properties of the object being edited. This is achieved by injecting the information from the input image into the logit space. Dotted red box indicates the masked region.
Discrete diffusion models are designed for inherently discrete data such as text or image tokens (esser2021taming, ). They adapt the diffusion framework to discrete spaces by defining appropriate transition kernels that corrupt and restore discrete data (hoogeboom2021argmax, ; austin2021structured, ; gu2022vector, ). Prominent examples include multinomial diffusion (hoogeboom2021argmax, ; gu2022vector, ), D3PM (austin2021structured, ), and masked generative models like MaskGIT (chang2022maskgit, ), Muse (chang2023muse, ). Despite their success in generation tasks, discrete diffusion models face limitations in controlled content editing. For instance, masked generative models achieve image editing through masked inpainting, where regions are masked and regenerated based on new conditions. However, this approach, as illustrated in Figure 1, lacks the ability to inject information from the masked area into the inpainting process, limiting fine-grained control over the editing outcome.
@@
Moreover, existing ODE-based inversion techniques developed for continuous diffusion models are not directly applicable to discrete diffusion models due to inherent differences in data representation and diffusion processes. This gap hinders the ability to perform precise inversion and controlled editing in discrete spaces. Thus, we propose DICE (Discrete Inversion for Controllable Editing), the first inversion algorithm for discrete diffusion models to the best of our knowledge. Our method extends the stochastic inversion approach to discrete diffusion models, including both multinomial diffusion and masked generative models. The core idea is to record the noise sequence to recover a stochastic trajectory in the reverse diffusion process. Specifically, given an artificial trajectory where latent states have low correlation, we fit reverse sampling steps to this trajectory and save the residuals between targets and predictions. This process imprints the information of the original input data into the recorded residuals. During editing or inference, the residuals are added back, allowing us to inject and control the amount of information introduced into the inference process.
Our approach enables accurate reconstruction of the original input data and facilitates controlled editing without the need for predefined masks or attention map manipulation. It provides a flexible framework for fine-grained content manipulation in discrete spaces, overcoming the limitations of existing methods. We validate the effectiveness of DICE through extensive experiments on both image and text modalities. We evaluate our method on models such as VQ-Diffusion (gu2022vector, ), Paella (rampas2022novel, ), and RoBERTa (liu2019roberta, ), demonstrating its versatility across different types of discrete generative models. Additionally, we introduce a novel text-editing dataset to further showcase our method’s capabilities and to facilitate future research in this area. Our contributions can be summarized as follows:
@@ 2 Related Work
Discrete diffusion. Generative modeling over discrete spaces using diffusion models has been extensively explored in recent years. Early developments include (sohl2015deep, ), Argmax Flows and Multinomial Diffusion (hoogeboom2021argmax, ), and D3PM (austin2021structured, ), which model the forward process as a discrete-time, discrete-state Markov chain. These methods train a neural network to reverse this process via a variational objective. Subsequent works such as (esser2021imagebart, ) and (gu2022vector, ) leveraged VQ-GAN to tokenize images, enabling efficient non-autoregressive generation strategies such as MaskGIT (chang2022maskgit, ), Muse (chang2023muse, ), and MMVID (han2022show, ) that iterate masking and prediction. Drawing inspiration from continuous diffusion models trained via score matching (song2019generative, ), recent methods introduce discrete analogues through ratio matching (lou2023discrete, ; meng2022concrete, ), which learn unnormalized density ratios. Discrete flow matching has also been proposed in this direction (gat2024discrete, ). In natural language processing, models such as BERT (devlin2018bert, ) and RoBERTa (liu2019roberta, ) have been interpreted as instances of discrete denoising (wang2019bert, ). More recently, there has been a surge of interest in developing diffusion-based language models (lou2023discrete, ; zheng2023reparameterized, ; shi2024simplified, ; sahoo2024simple, ; nie2025large, ; dream2025, ; arriola2025block, ), with successful systems demonstrating strong scalability and competitive performance with autoregressive LLMs.
Diffusion inversion. Diffusion inversion is the problem of taking an image and a text prompt that describes it and finding a noise latent that would generate the exact same image. More broadly, it refers to recovering the latent or noise vector that reproduces a given input under a diffusion model. Traditional approaches include deterministic inversion via neural ODEs chen2018neural , such as DDIM inversion song2021denoising and flow matching lipman2022flow ; liu2022flow , which reverse learned forward trajectories. Stochastic methods based on SDEs song2020score , including CycleDiffusion wu2022unifying and DDPM Inversion huberman2024edit , reconstruct the input by tracking noise along stochastic paths. To improve inversion quality, ReNoise garibi2024renoise ; pan2023effective applies fixed-point iterations, and GNRI samuel2023lightning uses a Newton-Raphson scheme to solve the DDIM inversion equation. Null-text Inversion mokady2022null improves reconstruction by optimizing null embeddings at test time, while Negative-Prompt Inversion miyake2023negative ; han2024proxedit introduces a closed-form approximation that reduces runtime without sacrificing fidelity. Our approach generalizes DDPM Inversion to discrete diffusion models, enabling effective inversion across both continuous and discrete domains.
Inversion-based image editing. A large body of diffusion-based image editing methods are grounded in DDIM inversion song2021denoising , which serves as the basis for reconstructing editable latents. These approaches often incorporate additional guidance mechanisms for controlled manipulation. For example, Prompt-to-Prompt hertz2022prompt modifies cross-attention maps, while Plug-and-Play tumanyan2023plug , TF-ICON lu2023tf , and StyleAligned hertz2024style expand this to self-attention layers. In contrast, DDPM inversion-based methods huberman2024edit provide user-friendly alternatives by avoiding complex attention map operations. They can be integrated with semantic guidance techniques such as SEGA brack2023sega , and this combination is exemplified by LEDITS tsaban2023ledits , which enables real image editing through DDPM inversion with semantic control. Other editing methods such as InstructPix2Pix brooks2023instructpix2pix rely on supervised fine-tuning over synthetic pairs and do not involve inversion, while Pix2PixZero parmar2023zero focuses on image-to-image translation using DDIM inversion with continuous diffusion.
@@ 3.1 Preliminaries
Masked generative modeling. Masked generative modeling is widely used in representation learning for both natural language processing and computer vision. It works by masking parts of the input and training the model to reconstruct the missing data. In models like BERT (devlin2018bert, ) and RoBERTa (liu2019roberta, ), masked tokens ([MASK]) are predicted based on the surrounding context, excelling in text completion and embedding representation learning. For image generation, Paella (rampas2022novel, ) adapts this approach for text-conditional image generation by renoising tokens instead of masking. The inference process in masked generative models typically involves iterative renoise/remask and repredict steps.
Multinomial Diffusion. Denoting $\bm{x}_{0}\in\{1,\ldots,K\}^{D}$ as a data point of dimension $D$. We use $\bm{v}({x}_{t}^{(i)})$ to denote the one hot column vector representation of the $i$-th entry of $\bm{x}_{t}$. To simplify notation, in the following we drop index $i$ and any function that operates on vector $\bm{x}_{t}$ is populated along its dimension. Diffusion model defines a markov chain $q(\bm{x}_{1:T}|\bm{x}_{0})=\Pi_{t=1}^{T}q(\bm{x}_{t}|\bm{x}_{t-1})$ that gradually adds noise to the data $\bm{x}_{0}$ for $T$ times so that $\bm{x}_{T}$ contains little to no information. Discrete diffusion model (hoogeboom2021argmax, ; austin2021structured, ; gu2022vector, ) proposed an alternative likelihood-based model for categorical data, and defines the forward process following:
@@
with $p_{\theta}(\tilde{x}_{0}|x_{t})$ is parameterized by a neural network. We gradually denoise from $x_{T}$ to $x_{0}$ using 3. For numerical stability, the implementation uses log space instead of probability space. Masked generative models can be viewed as a special case of multinomial diffusion models with an additional absorbing state (or the [MASK] state). Its training objective can be viewed as a reweighted ELBO (bond2022unleashing, ).
@@
16:  $\bm{x}_{t-1}\leftarrow\tilde{\bm{x}}_{0}\odot(\bm{1}-\bm{m}_{t-1})+\bm{n}\odot\bm{m}_{t-1}$
@@
2:for $t$ from $1$ to $\tau$ do
3:  $\bm{x}_{t}\sim q(\bm{x}_{t}|\bm{x}_{0})$ $\triangleright$ Independent q-sample using Eq 5
@@
6:for $t$ from $\tau$ to 11 do
7:  $\hat{\bm{y}}_{t-1}\leftarrow\log(\bm{\pi}_{\theta}(\bm{x}_{t},\bm{c},t))$ $\triangleright$ Log posterior using Eq 3
8:  $\bm{z}_{t}\leftarrow\bm{y}_{t-1}-\hat{\bm{y}}_{t-1}$ $\triangleright$ Eq 6
@@
3.2 Discrete Inversion for Controllable Editing
Non ODE-based inversion. ODE-based generative models, such as DDIM and flow matching, define an ODE trajectory. Due to the deterministic nature of ODEs, inversion can be achieved by solving the ODE using the Euler method in forward direction, ensuring reconstruction based on the inherent properties of the ODE. In contrast, another line of research focuses on SDE-based models, such as CycleDiffusion (wu2022unifying, ) and DDPM Inversion (huberman2024edit, ). Broadly speaking, these approaches ensure reconstruction by recording the noises or residuals that are required to reproduce the stochastic trajectory. CycleDiffusion records the Gaussian noise $\bm{z}_{t}$ during sampling from posterior $p(\bm{x}_{t-1}|\bm{x}_{t},\bm{x}_{0}=\bm{x}_{0})$ and injects information of the input signal by feeding the true $\bm{x}_{0}$. DDPM Inversion, on the other hand, incorporates information into $\bm{z}_{t}$ by fitting the reverse process into an artificial stochastic trajectory obtained by independent q-sample. For both CycleDiffusion and DDPM Inversion, the key idea is to utilize the Gaussian reparameterization trick, $x=\mu+\sigma z\Leftrightarrow x\sim\mathcal{N}(x;\mu,\sigma^{2})$, and keeping track of the “noise” that could have generated the sample from mean. For discrete diffusion models, we utilize the Gumbel-Max trick (maddison2014sampling, ; jang2016categorical, ), $x=\operatorname*{arg\,max}{\left(\log(\bm{\pi})+\bm{g}\right)}\Leftrightarrow x\sim\text{Cat}(x;\bm{\pi})$. Figure 2 provides an intuition of the proposed method.
Inverting masked generative models. In masked generative modeling, the stochastic trajectory $\bm{x}_{t}$} is constructed according to the specific inference algorithm of the model in use. For example, in Paella rampas2022novel , the masking is inclusive, meaning that as the time step $t$ increases, the set of masked tokens grows. In contrast, the Unleashing Transformer bond2022unleashing employs random masking at each step, where masks are generated independently using the q-sample function. Without loss of generality, we define a denoiser function $\mathcal{D}_{\theta}$ (parameterized by $\theta$). This denoiser outputs the logits of the predicted unmasked data given the noisy tokens $\bm{x}_{t}$. Unlike DDPM or multinomial diffusion, where $x_{t-1}$ is not sampled from a posterior given $x_{t}$, the inference of masked modeling takes a different approach. In masked modeling, $x_{t}$ is obtained from sampled $\hat{x}_{0|t}$ by re-noising. Since the categorical sampling occurs when drawing from the denoiser’s predicted logits, we accordingly define a corresponding latent sequence: